OpenStudy (anonymous):

Simplify the expression: cot x sin x - sin quantity pi divided by two minus x. + cos x

OpenStudy (ali2x2):

why dont you use wolfram or tiger-algebra

OpenStudy (ali2x2):

so easy too :P im not telling you you have to just a choice of action

OpenStudy (anonymous):

I could, but this helps teach me more.

OpenStudy (anonymous):

really need the equation tool for this almost impossible to read in english

OpenStudy (anonymous):

cot(x)sin(x)−sin((π2)−x)+cos(x)

OpenStudy (ali2x2):

ok

OpenStudy (ali2x2):

amazing how a mod agreed with me :3

Parth (parthkohli):

cot eks sin eks equals cos eks divided by sin eks times sin eks equals cos eks sin quantity pi divided by two minus eks equals cos eks so the expression boils down to cos eks minus cos eks plus cos eks

OpenStudy (anonymous):

Oh, and the answer choices are: cos(x) sin(x) 2sin(x) 2cos(x)

OpenStudy (ali2x2):

sin(x) + 2 cos(x)

OpenStudy (anonymous):

Yeah, @ParthKohli I got that part, but I can't figure out how to get one of the answer choices from it.

OpenStudy (anonymous):

Or at all, really....

OpenStudy (anonymous):

@ali2x2 That's not one of the possible answers.

OpenStudy (anonymous):

first off \[\sin(\frac{\pi}{2}-x)=\cos(x)\]

OpenStudy (ali2x2):

sorry that i couldnt help mabey @satellite73 will have better luck, MUCH BETTER LUCK

OpenStudy (ali2x2):

maybe

OpenStudy (anonymous):

@satellite73 That's an identity?

OpenStudy (anonymous):

@satellite73 ?

OpenStudy (anonymous):

Oh, whoops, @ParthKohli you solved it but it was confusing to look at, the final answer is cos(x).